OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -9 or x + 5 greater than or equal to 12

OpenStudy (anonymous):

a. A number line with a closed circle on -6, shading to the left, and an open circle on 7, shading to the right. b. A number line with an open circle on -6, a closed circle on 7, and shading in between. c. A number line with an open circle on -6, shading to the left, and a closed circle on 7, shading to the right. d. A number line with a closed circle on -6, an open circle on 7, and shading in between.

OpenStudy (texaschic101):

x - 3 < -9 x + 5 >= 12 x < -9 + 3 x >= 12 - 5 x < - 6 X >= 7 open circle on -6, shading to the left and closed circle on 7,shading to the right